OpenStudy (anonymous):

NEED HELP!! What is 1+2+3+4+5+6.......999

OpenStudy (anonymous):

do you just need the sum?

OpenStudy (anonymous):

do you mean 1+2+3.... all the way to 998+999??

mathslover (mathslover):

1+2+3+____+999 it is an arithmetic progression series

mathslover (mathslover):

common difference = a_2 - a_1 = 2-1 = 1

OpenStudy (anonymous):

go to wolframalpha.com and type in 1+2+3+4+5+6+...+999=

mathslover (mathslover):

so here i go with the formula .. : \[\large{s=\frac{n}{2}[2a+(n-1)d]}\]

mathslover (mathslover):

given a = 1 , n =999 \[\large{\frac{999}{2}[2+998]}\] \[\large{\frac{999}{2}[1000]}\]= 499500

mathslover (mathslover):

got it @vmathhelp ? ..

OpenStudy (anonymous):

Yes, and thank you

OpenStudy (anonymous):

Here is a nice way to figure it out: Say you add up the numbers 1 through 100 so you have 1+2+3+.....+98+99+100 So you add up the first and the last 1+100=101 add up 2+99=101 add up 3+98=101 you get the point, you have 50 pairs of these right? so you have 50(101)=5,050 and i found this online: http://mathforum.org/library/drmath/view/57919.html

OpenStudy (anonymous):

isn't it 500,499

mathslover (mathslover):

no problem ..

mathslover (mathslover):

no @vmathhelp it will be 499,500

mathslover (mathslover):

999 * 500 = 499500

OpenStudy (anonymous):

I didn't post answer i posted the logic behind it

OpenStudy (anonymous):

500,500

mathslover (mathslover):

notice: +100....

mathslover (mathslover):

your series dont contain 100 and so on

OpenStudy (anonymous):

formulas aren't helpful if you don't understand their derivation (where they're coming from) i hardly memorized any formulas during my 4 years of studying it at university level and did splendid b/c of understanding derivations...

mathslover (mathslover):

here is the proof @mashe

OpenStudy (anonymous):

\[\sum_{k=1}^{k=n} k = n (n-1)/2\]